Greatest common factor?
24xy^3+42xy^2+54x^4y^4

Answer by greenestamps(13198)



Constant: 24, 42, 54 ==> GCF = 6
variable x: x, x, x^4 ==> GCF = x
variable y: y^3, y^2, y^4 ==> GCF = y^2

GCF of the whole expression: 6xy^2